1.Find the Maclaurin series for thefunction f(x)= (x^2+ 2-x)5/2and its radiusofconvergence


1
Expert's answer
2022-05-06T13:02:00-0400

Remind that Maclaurin series is the Taylor series considered at point "x=0". The Maclaurin series has the form: "f(0)+f'(0)x+f''(0)\frac{x^2}{2!}+f'''(0)\frac{x^3}{3!}+...". Consider the function "g=(2-z)^{\frac{5}2}=2^{\frac52}(1-\frac{z}2)^{\frac52}". Maclaurin series for the function "(1-x)^{\alpha}" has the form: "(1-x)^{\alpha}=\sum_{n=0}^{+\infty}C_{\alpha}^n(-1)^nx^{n}," where "C_{\alpha}^n=\frac{\alpha(\alpha-1)...(\alpha-n+1)}{n!}". It is a known series and it converges for "|x|<1". We get: "g=(2-z)^{\frac{5}2}=2^{\frac52}(1-\frac{z}2)^{\frac52}=\sum_{n=0}^{+\infty}C_{\alpha}^n(-1)^n\left(\frac12\right)^nz^{n}". Its radius of convergence is: "|z|<2". Now we set "z=x-x^2". We receive: "\sum_{n=0}^{+\infty}C_{\alpha}^n(-1)^n\left(\frac12\right)^n(x-x^2)^{n}".

It remains to expand it using the binomial formula "(x-x^2)^n=\sum_{k=0}^nC_n^kx^{k}(-1)^{n-k}x^{2n-2k}",where "C_n^k=\frac{n!}{k!(n-k)!}". We get: "\sum_{n=0}^{+\infty}C_{\alpha}^n(-1)^n\left(\frac12\right)^n\sum_{k=0}^nC_n^k(-1)^{n-k}x^{2n-k}". The series converges for "|x^2-x|<2". It is enough to solve the latter inequality to obtain the radius of convergence. Consider the expression: "x(x-1)". It is positive for "x\in(-\infty,0)\cup(1,+\infty)". For "x\in(-\infty,0)\cup(1,+\infty)" we receive: "x^2-x-2<0". The latter is equivalent to: "(x+1)(x-2)<0". The solution is "(-1,0)\cup(1,2)". For "x\in[0,1]" we get: "x^2-x+2>0". The latter is satisfied for "x\in[0,1]." Thus, the Maclaurin series has the form: "\sum_{n=0}^{+\infty}\sum_{k=0}^nC_{\alpha}^nC_n^k(-1)^{2n-k}\left(\frac12\right)^nx^{2n-k}" and it converges for "x\in(-1,2)".

Answer: the Maclaurin series is: "\sum_{n=0}^{+\infty}\sum_{k=0}^nC_{\alpha}^nC_n^k(-1)^{2n-k}\left(\frac12\right)^nx^{2n-k}". It converges for "x\in(-1,2)."